OpenStudy (anonymous):

Medal if you could help and explain:) During a clothing stores sale , the regular shirts are discounted by $5.00.There is a state sales tax of 5% and the $5.00 discount is applied before the sales tax is calculated. a. Write an expression the regular price r of a t shirt with the $5.00 discount applied. b.Write a rule for the funtion p (r) that expresses the final price p of a t shirt with the discount applied and sales tax added. c. How much would you pay for a t shirt on the sales day priced at $15.50?

OpenStudy (anonymous):

The answer is around $11.02

OpenStudy (anonymous):

a. The expression for the regular price r with $5.00 discount is ...$10.50 b. with $5.00 discount sales tax..52 c..$11.02

OpenStudy (anonymous):

@shamil98

OpenStudy (lasttccasey):

So for part a, it's asking for an expression: `r - 5` where r is the regular price. For b, it wants you to make this a function and incorporate the tax: `p(r) = (r - 5)*(1 + tax)` For c, you're simply inserting the numbers for r = 15.50 and tax = 0.05

OpenStudy (anonymous):

@Garadon_Shagan

OpenStudy (anonymous):

Ok so do you get $11.02 for the c. part ?

OpenStudy (lasttccasey):

Well if you plug it in you'll find: `p(15.50) = (15.50 - 5)*(1 + 0.05)` = `(10.5) * (1.05)` = `$11.025`

OpenStudy (anonymous):

Awesome see I know the answer but have a problem putting it into form .Thank you for ur help ,!!